OpenStudy (anonymous):

I Need the answer pleaseee. Below are the steps to solve an equation: Step 1: |x − 4| + 1 = 8 Step 2: |x − 4| = 8 − 1 Step 3: |x − 4| = 7 Which of the following is a correct next step to solve the equation? x − 4 = −7 x + 4 = −7 −x − 4 = 7 x + 4 = 7

OpenStudy (welshfella):

the expression that's in the vertical lines can be negative or positive, The vertical lines mean the absolute value ( always positive)

OpenStudy (welshfella):

so we can write x - 4 = 7 or x -4 = -7

OpenStudy (welshfella):

so which of the 3 options do you think is correct?

OpenStudy (welshfella):

* 4 options

OpenStudy (anonymous):

so the answer is A?

OpenStudy (anonymous):

i meant B, sorry

OpenStudy (welshfella):

one of the equations in my second post is also one of the options

OpenStudy (welshfella):

you were right first time it A x - 4 = -7

OpenStudy (welshfella):

it cant be x + 4 as its x - 4 in the brackets

OpenStudy (anonymous):

okay thanks

OpenStudy (welshfella):

yw